On Jan 2, 2020 Albert Aloysius Zander Equipment, Inc. (AAZEQ) and Mole Excavating (MOLEX) entered into a four-year lease contract for an IVAN-84D, which AAZEQ also offered for sale at $110,000. AAZEQ estimated that each machine could either be refurbished at the end of the lease or scrapped for parts, and that the Residual value of the equipment would thus be $18,000. The economic life of the asset was six years. MOLEX incurred $4,400 in related direct costs for the lease. The lease required payments at the end of each year, commencing Dec 31, 2020. AAZEQ used a 4.7% discount rate and MOLEX’s IBR was 7.2%.

Required: Supply the amounts requested on the answer sheet. If the item description is not applicable to the facts of the problem, put NA in the box.

Implicit rate on the lease:

Lease receivable:

Lease expense:

Lessor gross profit:

Lease liability:

Right of use asset:

Deferred gross profit:
Answer #1

MOLEX is lessor of the contract and accordingly following are Calculations

Implicit rate of return : 7.2%

accordingly PV Factors are as follows:

0.93, 0.87, 0.81, 0.76

Calculation of Annual Installment receivable :

Annual Installment=Cost of Asset / PV

Cost of Asset = 110000+4400= 114400

(Expenses directly Incurred on lease are capitalised)

= 114400/ 3.37

$ 33,927

Calculation of Depreciation = (114400-18000)/4

=24100

Year I Year II Year III Year IV

Lease Receivable 33927 33927 33927 33927

Lease Expenses 24100 24100 24100 24100

Gross Profit 9827 9827 9827 9827

PV 0.93 0.87 0.81 0.76

Present Value 9166 8551 7976 7440

AAZEQ is Lessee of the contract

Following are some Calcuations of Lease Contract

Discount Rate = 4.7%

PV ( 0.96, 0.91, 0.87, 0.83)

Lease Liability will be in books of lessee = PV (Minimum Lease Payment + Guranteed Residual Value)

Annual Installment = Cost of Asset/ PVF

= 110000/ 3.57

= 30805

Particulars Year I Year II Year III Year IV

Annual Installment 30805 30805 30805   30805

PV Factor 0.96 0.91 0.87 0.83

Present Value 29573 28033 26801 25568

Guaranteed Value 18000

PV 0.83

Present Value 14,940

Total 124915 (109975+14940)

Deferred Gross Profit

23,969 ( YearII 8551+Year III 7976+Year III 7440)